A Complete Bibliography of Publications in the journal Groups Complexity Cryptology

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 November 2019 Version 1.05

Title word cross-reference

{+, −, ×, ÷, &} [Sch16]. 1 [ET16]. 4 [Ore15]. 5 [EE17]. 5 × 5 [HHMHM15]. Aut(F_{n}) [Ibr12]. F [ERW12, ET16, TY15, Was11]. F_{m} × Z^{n} [Sah15]. ⟨m, m, m⟩ [HHMHM15]. Z_{q} [AGG10]. Zn⋆ [Ano18a]. O((log q)^{2}) [Sch16]. n [RS14]. p [Ano17a]. p^{n} [Iva15]. R [MZ09]. BS(1, 2) [Mil14].

-braids [Ore15]. -counter [ET16]. -groups [Ano17a]. -powered [MZ09].

2-dimensional [Asb18].

AAG [Rom19]. abelian [Ano17a, BMV15, Cle09, Dea09, Eic16, KA09]. action [Ibr12].

Actions [BD09, KMS14]. active [Neu12]. additive [Cle09]. admitting [GLS15]. agreement [CM15, GG19, MU09, Ush16]. algebra [Bre11]. Algebraic [BBC^{+}14, JK10, Kre09, She10, BD09, BD11, Ore15, Rom18, She16, She19, AAGG16]. algorithm [Bac15, CL13, CJ12, HHMHM15, MU14, STT11]. Algorithmic [Ore15, MU11b]. Algorithms [Loh12, BM11, HM12, Wil12]. Almost [GLS09]. alternative [gRHR09]. amalgams [FMR09]. analogs [BMV15]. Analysis [KPU18, GLS15]. Andrews [PU19]. Anshel [MU09, MU09]. apartments [CL13]. application [GLS15, HHMHM15]. applications [CJ12, KMK19, SS16]. approach [CR16]. arithmetic [KMK19, Mac09]. Artin [HR13, Kra13]. aspects [FGRS14]. associative [Bre11, KT13]. assumption [Ano13, Ano18a, Ano19a]. assumptions [Neu12]. asymmetric [Kra13]. asymptotic [CFR10]. asymptotics [Wil12]. Ate [FC16]. attack [LU09, MR15, Rom16]. attacks [JK10, Kre09]. Authenticated [Ush16]. Authentication [GS09, GN12, GS12, LU09]. automatic [ET16, Gil14]. automaticity [HR13]. automaton [GG19]. automorphism [Eic16]. automorphisms [She19].

base [MR14]. based [AAGG16, FMR13, GG19, GK16, KPU18, LU09, MZJ17, Rom15, Shp14, TY15]. Baumslag [Cle09]. Ben [Ros09]. bilinear [Das15]. biography [Ros09]. biometric [GN12]. bits [AK16]. BN [DG16]. Boolean [Ryb17]. boundary [Bac15]. braid [BD09, Ito10, KLT09, LU09, SW19]. braids [Ore15].

cancellation [Gil14]. canonical [KMT19]. case [Mya09, NR16]. Cayley [AGG10, Ant11, GM18]. certain [Ano18a, LU09]. Certifying [KN18]. Challenge [BBFT10]. cheater [PA19]. cheaters [PA19]. ciphers [BBC^{+}14]. circulant [Mah10]. class [AAGG16, BD11, Bac15, KS19]. classification [She10]. closest [FMR13]. Closure [KS19]. cogrowth [ERW12]. combinatorial [Bac15, BBFT10, Rom17]. commutative [KK12, MZ09]. commutator [Ush16]. compactness [SS17]. complete [GHP09]. complexity [LT09, Mya09, NR16, Ryb13, Ryb15, TY15]. component [Bac15]. Compositions [SS16]. compressed [Loh12]. computable [Mil11]. computation [DG16, FC16]. computational [Ano13, Ano17a, Ano19a]. compute [Bac15, Bre11]. Computing [Sch16]. conditions [SS17]. conjecture [FGRS18, PU19]. Conjugacy [PU19, BN16, CJ12, DDM12, KLT09, KLT10, LU09, MV17, Shn09, Vas11]. conjugation [GS09]. Constructing [Ano13, Ano19a]. construction [KMT19]. context [Bro14, Ho18, KS19]. context-free [Ho18, KS19]. Continuous [GN12]. convex [Sah15]. coordinate [She10]. coset [CL13]. counter [ET16]. covering [KNPS16]. covers [KMT19]. Cramer [KA09]. Cryptanalysis [MN15, Mon19, MU09, Rom17]. cryptographic [Rom19]. Cryptography [CM17, CJ12, GK16, KMK19, MR14, Rom18, SZ09]. cryptosystem [FMR13, GHP09, KA09, MZJ17, Rom15, Rom17]. Curtis [PU19]. curves [DG16, FC16, Fri17]. Cutting [Krö10]. Cyclic [DDM12, But17]. Cylinders [Ibr12].

decidability [Ven14]. Decision [KA09, SZ09]. decomposition [Bre11, MR15, Rom16]. decompositions [Wil12]. Decoy [Shp14]. Decoy-based [Shp14]. Dehornoy [BD09]. densities [CFR10]. diagonalizable [MZJ17]. diagrams [KMT19, MU11b]. diameter [AGG10]. differentiable [Iva15]. Diffie [Eft12]. diffusion [SDR19]. digital [KK12, STT11]. dimension [HK13]. dimensional [Asb18, Bre11]. Diophantine [GMO17, MR14, Ryb13]. Direct [SS17, Wil12]. discrete [Mah10, MU14, Sch16]. distributive [KT13]. doubles [Pue16].

each [Asb18]. echelon [Ros13]. Effective [KMT19]. elementary [Ano17a, FGRS18, Fri17, GLS15]. elements [Asb18, But17]. elliptic [FC16, Fri17]. embedding [Cho09]. encryption [GKS18b, Rom15]. endomorphism [FFK11]. Equations [Rom12, CR16, GMR11, KMT19, Tim17]. eraser^{T M} [AAGG16]. escrow [Das15]. establishment [KT13, Neu12]. estimates [KN18]. Evolutionary [CJ12, CR16]. examples [SW19]. exchange [Eft12, KKS13]. Existence [Mac09, Wil12]. expanding [LW12]. exponentiation [DG16]. exponents [GLS15]. extensions [DE15].

factor [Cle09]. Factoring [AK16, Ano13, Ano18a, Ano19a]. Faithful [FR13, FR11]. families [Ano17a, LW12]. family [Ano13, Ano18a, Ano19a]. fast [HHMHM15]. Faster [FC16]. fields [Mil11]. final [DG16]. find [CL13]. Fine [Ros09]. finite [Ano13, Ano17a, Ano19a, AT15, Bre11, Ito10, MU14, Pue16]. finite-dimensional [Bre11]. finitely [CFR10, Chi14, Eic16, HK13, SZ18]. fixed [BM11]. Free [But17, Ano13, Ano17a, Ano18a, Ano19a, Ant11, Bro14, Cle09, Das15, Dea09, EE17, GLS09, GMR11, HK13, Ho18, KMS14, KMT19, KS19, MR14, Ros13, SZ18, Tim17, Vas11]. Friends [Mil14]. Front [Ano17b, Ano18b]. Frontmatter [Ano17c, Ano18c, Ano19b, Ano19c, Ano19d]. Fuchsian [Mac09]. fully [GKS18b]. function [GM18, Mon19]. functions [AAGG16, GN12, GS13, Iva15, Mya09, SS16].

galore [Kre09]. gaps [KN18]. Garside [SW19, KLT10]. general [Ano13, Ano18a, Ano19a, Rom18]. generalisation [Kra13]. Generalized [Gil14, BBC^{+}14, BN16, gRHR09]. generated [CFR10, Eic16, HK13, SZ18]. generating [BD11]. Generic [FMR09, Mya09, NR16, Ryb13, Ryb17, CFR10, Ryb15, STT11]. geodesic [ER10, HR13]. geometries [CL13]. geometry [She10, She19]. Goldfeld [MU09]. graph [AGG10, ET16, NR16, Ryb15]. graphs [Ant11, AT15, Krö10, MU11a]. Grigorchuk [MV17]. Group [DE15, Ven14, Asb18, BD09, BBFT10, Cle09, Dea09, Eic16, ERW12, ET16, FFK11, FMR09, Fri17, GLS15, GK16, KKS13, LU09, Mah10, MV17, MN15, MU14, Pue16, She19, Shp10, TY15, Was11]. group-based [GK16]. Group-theoretic [Ven14]. Groups [Bro14, GKS18a, Ano13, Ano17a, Ano19a, Ant11, Asb18, BD11, Bac15, BN16, But17, CFR10, Chi14, Cho09, Cle09, Con10, CJ12, CR16, EE17, ER10, FR11, FR13, FGRS14, FGRS18, GLS09, GMO17, GMR11, Gil14, GG19, HK13, HR13, Ito10, KMK19, KLT09, KLT10, KNPS16, KMT19, KS19, LT09, LW12, Mac09, MZ09, MZJ17, Mil11, MT17, MU11a, MU11b, MR14, NBR10, Osi09, RS14, Rom12, Ros13, SW19, Shn09, SZ18, Tim17, Vas11, Wei10, gRHR09]. Growth [FFK11].

hard [GN12]. hard-to-invert [GN12].

hardness [Ryb17]. hash

[AAGG16, GM18, Mon19]. hashing [SS16]. Heisenberg [BN16]. Hellman [Eft12]. Hessian [CM17]. Hirsch [EE17]. Holmes [GS12]. Hom [KMT19]. Hom-diagrams [KMT19]. homology [AT15, NBR10]. homomorphic [GKS18b]. hulls [Sah15]. Hurwitz [Con10]. Hydra [Pue16]. hyperbolic [KMT19, NBR10, Osi09].

identification [PA19]. II [FR13].

improved [Rom19]. indistinguishable [LW12]. induced [Ibr12]. inert [Ros13]. Infinite [KMS14, FGRS14]. information [Shp14]. integer [Ano13, Ano18a, Ano19a]. integral [GLS15]. intersection [Ros13]. intractability [Ano13, Ano18a, Ano19a]. introducing [Neu12]. introduction [Mil11]. invariant [BM11]. invariants [MZJ17]. invert [GN12]. irreducible [She16]. isomorphic [Was11]. Isomorphism [LW12, EE17, NR16, Ryb15].

Kampen [MU11b]. Kaplansky [FGRS18].

Key [CM15, Das15, GG19, Eft12, FMR13, GKS18b, GHP09, KA09, KKS13, KT13, MZJ17, MU09, Neu12, Rom15, Rom17, SZ09, Ush16]. Key-escrow [Das15]. Knapsack [MT17].

lacunary [Osi09]. language [Ho18, TY15]. large [MU11a, Wei10]. Larue [BD09]. latin [CZ10]. lattice [BMV15, CCR11]. law [Fri17, MU11a]. layers [SDR19]. leak [GS12]. least [AK16]. left [KT13]. Lemieux [MU09]. length [EE17, Ore15]. like [BBC^{+}14]. limit [FR11, FR13]. linear [Asb18, But17, MR15, SS16]. linearly [She16]. locally [GLS09]. Log [MV17]. Log-space [MV17]. logarithm [Mah10, MU14]. logarithms [Sch16]. look [KM13].

Magnus [GLS09]. mapping [BD11, Bac15]. mapping-class [BD11]. matrices [Eft12, KKS13, Mah10, MN15, MU14]. matrix [GS09, HK13, HHMHM15, Kas10]. matter [Ano17b, Ano18b]. max [RS14]. max- [RS14]. maximal [Mac09]. McCool [BD11]. MDS [SDR19]. Memory [DG16]. Memory-saving [DG16]. Metabelian [Dea09, GLS15, MR14, RS14, Tim17]. method [BD11, GS12]. model [FC16, Mil11]. modification [FMR13]. modular [Mon19]. moduli [AK16]. modulo [Iva15]. monoids [Cho09, Kra13, She10]. most [AK16, EE17]. multi [AK16, Das15, Ibr12]. multi-cylinders [Ibr12]. multi-power [AK16]. multi-signature [Das15]. multiple [CJ12, Ho18, KS19]. multiplication [HHMHM15].

n [Ho18]. natural [She10]. Nielsen [KPU18]. nilpotent [Dea09, EE17, GMO17, HK13, MZ09, MT17]. No [GS12]. No-leak [GS12]. Non [BMV15, KK12, KT13, KA09, KM13, Mac09, Mah10].

Non-abelian [BMV15, KA09]. Non-associative [KT13]. Non-commutative [KK12]. non-existence [Mac09]. non-singular [Mah10]. non-uniformity [KM13]. noncommutative [Eft12]. nonlinear [Rom16, SDR19]. note [FR11, KLT09, NBR10]. number [KNPS16]. numbers [Asb18, Cle09, MU11a, She10]. numerical [KN18].

one [Bac15, GS13, Mya09]. one-way [GS13, Mya09]. open [GK16]. operations [Sch16]. orbit [Ven14]. order [Asb18]. Orderable [FGRS18]. ordered [She16]. orderings [Ito10]. orientable [Bac15]. oriented [SDR19].

pairing [DG16, FC16]. pairings [Das15]. Palindromic [RS14]. parallel [CR16]. parallels [CM15]. party [Neu12]. passive [Neu12]. password [BBFT10]. pencil [CM17]. points [BM11]. poly [Bro14]. poly-context-free [Bro14]. polycyclic [CR16, GMO17, GK16]. Polynomial [Vas11, GLS15]. Power [MZ09, AK16]. Power-commutative [MZ09]. powered [MZ09]. Practical [GKS18b, LU09]. presence [PA19]. Presentations [Kas10, Bac15, GMO17, Gil14]. presented [Chi14]. primes [AK16]. private [FMR13, GKS18b]. private-key [GKS18b]. probabilistic [Rom15]. problem [BN16, Bro14, CJ12, EE17, Ho18, KLT09, KLT10, Mah10, MV17, MT17, Mon19, MU14, NR16, PU19, Ryb13, Ryb15, Ryb17, Shn09]. problems [DDM12, ER10, GMO17, GKS18a, GK16, LT09, MU11b, SZ09, Shp10]. product [Dea09, HHMHM15, HM12, Wil12]. products [RS14, SS17, Vas11]. projective [Asb18]. proof [Fri17, Krö10]. proofs [BD09]. properties [BBC^{+}14, CFR10, DE15, KS19]. property [HHMHM15, HM12]. protocol [Eft12, LU09, MU09, Rom19, Ush16]. Pseudo [Ano17a, Ano13, Ano18a, Ano19a]. Pseudo-free [Ano17a, Ano13, Ano18a, Ano19a]. PTIME [BN16]. Public [KKS13, MZJ17, GHP09, KA09, Rom15, Rom17, SZ09]. Public-key [MZJ17, GHP09, Rom15].

Quantum [BM11, MU14]. quasiconvex [Sah15]. quasipositive [Ore15]. questions [GLS09].

R. [Was11]. Ramp [PA19]. Random [GMO17, GMR11, MU11b, AGG10]. Randomized [SDR19]. rate [FFK11]. rational [Cle09]. real [GLS15]. Recognition [Asb18, Ore15]. Reflections [FGRS14]. regularity [HR13]. relatives [Mil14]. remark [Tim17]. representations [FR11, FR13, HK13]. residually [Pue16]. response [BBFT10]. restricted [BN16, But17]. results [BD09]. revisited [Krö10]. Rewriting [Cho09, DDM12]. Rijndael [BBC^{+}14]. Rijndael-like [BBC^{+}14]. rings [Eft12, GKS18b, KKS13, Kas10, MN15, MU14]. rounding [BMV15]. RSA [AK16, Rom15]. rushing [PA19].

SAT [JK10]. SAT-solvers [JK10]. satisfiability [Ryb17]. saving [DG16]. scheme [Das15, FMR13]. schemes [CZ10, KPU18, Rom18]. Search [HM12, Shp10, CJ12, HHMHM15, KA09, PU19]. searching [Ryb15]. Secrecy [GS13]. secret [CZ10, FMR13, KPU18, PA19]. secure [GM18]. security [BBFT10, Neu12, Shp14]. Selmer [FC16]. semilattices [She16]. semilinear [GKS18a]. set [Asb18]. sets [She16]. sharing [AK16, CZ10, FMR13, KPU18, PA19]. Shephard [Wei10]. Sherlock [GS12]. shifted [KLT09, LU09]. short [Krö10, Ros09]. Shortlex [HR13]. Shoup [KA09]. shuffle [Shn09]. signature [Das15, STT11]. signatures [KK12]. significant [AK16]. simple [KNPS16]. singular [Mah10]. SLP [Loh12]. SLP-compressed [Loh12]. small [Gil14, KNPS16]. software [SDR19]. software-oriented [SDR19]. Solitar [Cle09]. solution [BN16, CJ12]. solvable [RS14, Vas11]. solvers [JK10]. solving [CR16]. Some [ER10, KMK19, SW19, FGRS14, GLS09, KNPS16]. Space [LT09, MV17]. special [DE15]. spectral [KN18]. spherical [Tim17]. sporadic [KNPS16]. squares [CZ10]. Stallings [Krö10]. status [GK16]. strings [Loh12]. Strong [MU11a]. structure [Krö10]. Subgroup [KLT10]. Subgroups [Was11, Ano18a, BM11, CCR11, Cle09, FMR09, KLT10, Ros13, Sah15]. submonoids [SZ18]. subset [Mon19]. subsurfaces [SW19]. sum [Mon19]. surfaces [Bac15]. survey [GK16, Loh12]. symmetric [KNPS16]. Symmetries [AT15]. system [MN15]. systems [Cho09, CR16, KT13].

test [HM12]. tetrahedron [gRHR09]. theorem [FMR13, GLS09, Krö10]. theoretic [Ven14]. theoretical [MR14]. theory [BBFT10, FGRS18, Mil11, Osi09, SW19, Shp10]. Thompson [ERW12, ET16, TY15, Was11]. three [Ore15]. Thurston [Ito10]. Thurston-type [Ito10]. time [Vas11]. Tits [gRHR09]. Torsion [Cle09, Chi14, EE17, HK13, KMT19, Mac09, Osi09]. Torsion-free [Cle09, KMT19]. transformations [KPU18]. transitive [Iva15]. Tree [CCR11, TY15]. Tree-based [TY15]. triangles [Wei10]. triple [HHMHM15, HM12]. triples [HHMHM15, HM12]. tropical [CM15, CM17]. Tsaranov [gRHR09]. Two [Neu12, Rom18]. Two-party [Neu12]. type [Ito10]. uniformity [KM13]. unipotent [But17]. universal [KMS14, Osi09, She19]. update [Con10]. Using [SZ09, BBFT10, Das15, Eft12, JK10, KKS13, MN15, Sch16].

varieties [SS17]. vector [FMR13]. version [GM18, Rom19]. virtually [Ant11, Eic16, SZ18].

way [GS13, Mya09]. weakly [Ano18a]. Wedderburn [Bre11]. whose [GKS18a]. width [RS14]. without [GS13, Neu12]. witness [Shp10]. word [Bro14, GKS18a, Ho18, LT09, Shn09]. words [KMS14]. wreath [RS14, Vas11].

Z [Ho18]. Zieschang [BD11].

References

Anshel:2016:CHF

[AAGG16] Iris Anshel, Derek Atkins, Dorian Goldfeld, and Paul E. Gunnells. A class of hash functions based on the algebraic eraser^{T M}. Groups. Complexity. Cryptology, 8(1):1–7, 2016. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Amir:2010:DRC

[AGG10] Gideon Amir and Ori Gurel-Gurevich. The diameter of a random Cayley graph of Z_{q}. Groups. Complexity. Cryptology, 2(1):59–65, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Akchiche:2016:FMP

[AK16] Omar Akchiche and Omar Khadir. Factoring multi-power RSA moduli with primes sharing least or most significant bits. Groups. Complexity. Cryptology, 8(1):47–54, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Anokhin:2013:CPF

[Ano13] Mikhail Anokhin. Constructing a pseudo-free family of finite computational groups under the general integer factoring intractability assumption. Groups. Complexity. Cryptology, 5(1):53–74, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Anokhin:2017:PFF

[Ano17a]  Mikhail Anokhin.  Pseudo-

free families of finite computational elementary abelian pgroups. Groups. Complexity. Cryptology, 9(1):1–??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

//www.degruyter.com/view/
j/gcc.2017.9.issue-1/gcc-
2017-0001/gcc-2017-0001.xml.

Anonymous:2017:FM

[Ano17b] Anonymous. Front matter. Groups. Complexity. Cryptology, 9(2):i, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Anonymous:2017:F

[Ano17c]  Anonymous.  Frontmatter.

Groups. Complexity. Cryptology, 9(1):i–??, May 2017. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2017.9.issue-1/gcc-2017-frontmatter1/
gcc-2017-frontmatter1.xml.

Anokhin:2018:CFS

[Ano18a]  Mikhail Anokhin.  A cer-

tain family of subgroups of Zn⋆ is weakly pseudo-free under the general integer factoring intractability assumption. Groups. Complexity. Cryptology, 10(2):99–??, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Anonymous:2018:FM

[Ano18b] Anonymous. Front matter.

Groups. Complexity. Cryptology, 10(2):i–iv, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Anonymous:2018:F

[Ano18c]  Anonymous.  Frontmatter.

Groups. Complexity. Cryptology, 10(1):i–??, May 2018. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2018.10.issue-1/gcc-2018-
frontmatter1/gcc-2018-frontmatter1.
xml.

Anokhin:2019:CPF

[Ano19a] Mikhail Anokhin. Constructing a pseudo-free family of finite computational groups under the general integer factoring intractability assumption. Groups. Complexity. Cryptology, 11(2):133–134, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Anonymous:2019:Fa

[Ano19b]  Anonymous.  Frontmatter.

Groups. Complexity. Cryptology, 11(1):i–??, May 2019. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

                                        tronic). URL https://www.
                                        degruyter.com/view/j/gcc.
                                        2019.11.issue-1/gcc-2019-
                                        frontmatter1/gcc-2019-frontmatter1.
                                        xml.

Anonymous:2019:Fb

[Ano19c]  Anonymous.  Frontmatter.

Groups. Complexity. Cryptology, 11(2):i–iv, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Anonymous:2019:Fc

[Ano19d]  Anonymous.  Frontmatter.

Groups. Complexity. Cryptology, 11(2):i–iv, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Antolin:2011:CGV

[Ant11] Yago Antolín. On Cayley graphs of virtually free groups. Groups. Complexity. Cryptology, 3(2): 301–327, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Asboei:2018:RDP

[Asb18] Alireza Khalili Asboei. Recognition of 2-dimensional projective linear groups by the group order and the set of numbers of its elements of each order. Groups. Complexity. Cryptology, 10(2):111–??, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Atchison:2015:SFG

[AT15] Benjamin Atchison and Edward C. Turner. Symmetries of finite graphs and homology. Groups. Complexity. Cryptology, 7(1):11–30, 2015. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Bacardit:2015:CAC

[Bac15] Lluís Bacardit. A combinatorial algorithm to compute presentations of mapping class groups of orientable surfaces with one boundary component. Groups. Complexity. Cryptology, 7(2): 95–115, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Babinkostova:2014:APG

[BBC^{+}14] Liljana Babinkostova, Kevin W.

Bombardier, Matthew C. Cole, Thomas A. Morrell, and Cory B. Scott. Algebraic properties of generalized Rijndaellike ciphers. Groups. Complexity. Cryptology, 6(1):37–54, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Baumslag:2010:CRP

[BBFT10] Gilbert Baumslag, Yegor Bryukhov, Benjamin Fine, and Douglas Troeger. Challenge response password security using combinatorial group theory. Groups.

Complexity. Cryptology, 2(1): 67–81, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Bacardit:2009:ABG

[BD09] Lluís Bacardit and Warren Dicks. Actions of the braid group, and new algebraic proofs of results of Dehornoy and Larue. Groups. Complexity. Cryptology, 1(1):77–129, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Bacardit:2011:ZMM

[BD11] Lluís Bacardit and Warren Dicks. The Zieschang–McCool method for generating algebraic mapping-class groups. Groups. Complexity. Cryptology, 3(2): 187–220, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Bonanome:2011:QAF

[BM11] Marianna Bonanome and Stephen Majewicz. Quantum algorithms for fixed points and invariant subgroups. Groups. Complexity. Cryptology, 3(2):329–348, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Begelfor:2015:NAA

[BMV15] Evgeni Begelfor, Stephen D. Miller, and Ramarathnam Venkatesan. Non-abelian analogs of lattice rounding. Groups. Complexity. Cryptology, 7(2):117–133, 2015. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Blaney:2016:PSR

[BN16] Kenneth R. Blaney and Andrey Nikolaev. A PTIME solution to the restricted conjugacy problem in generalized Heisenberg groups. Groups. Complexity. Cryptology, 8(1):69–74, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Bremner:2011:HCW

[Bre11] Murray R. Bremner. How to compute the Wedderburn decomposition of a finitedimensional associative algebra. Groups. Complexity. Cryptology, 3(1):47–66, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Brough:2014:GPC

[Bro14] Tara Brough. Groups with poly-context-free word problem. Groups. Complexity. Cryptology, 6(1):9–29, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Button:2017:FCG

[But17] Jack O. Button. Free by cyclic groups and linear groups with restricted unipotent elements. Groups. Complexity. Cryptology, 9(2):137–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Carbone:2011:TLS

[CCR11] Lisa Carbone, Leigh Cobbs, and Gabriel Rosenberg. Tree lattice subgroups. Groups. Complexity. Cryptology, 3(1):1–23, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Carstensen:2010:ADG

[CFR10] Celine Carstensen, Benjamin Fine, and Gerhard Rosenberger. On asymptotic densities and generic properties in finitely generated groups. Groups. Complexity. Cryptology, 2(2): 113–121, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Chiodo:2014:TFP

[Chi14] Maurice Chiodo. On torsion in finitely presented groups. Groups. Complexity. Cryptology, 6(1):1–8, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Chouraqui:2009:RSE

[Cho09] Fabienne Chouraqui. Rewriting systems and embedding of monoids in groups. Groups. Complexity. Cryptology, 1(1): 131–140, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Craven:2012:EAS

[CJ12] Matthew J. Craven and Henri C. Jimbo. Evolutionary algorithm solution of the multiple conjugacy search problem in groups, and its applications to cryptography. Groups. Complexity. Cryptology, 4(1):135–165, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Connor:2013:NAF

[CL13] Thomas Connor and Dimitri Leemans. A new algorithm to find apartments in coset geometries. Groups. Complexity. Cryptology, 5(1):75–89, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Clement:2009:TFA

[Cle09] Anthony E. Clement. Torsionfree abelian factor groups of the Baumslag–Solitar groups and subgroups of the additive group of the rational numbers. Groups. Complexity. Cryptology, 1(2):165–168, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Chauvet:2015:KAU

[CM15] Jean-Marie Chauvet and Eric Mahé. Key agreement under tropical parallels. Groups. Complexity. Cryptology, 7(2): 195–198, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Chauvet:2017:CTH

[CM17] Jean-Marie Chauvet and Eric Mahé. Cryptography from the tropical Hessian pencil. Groups. Complexity. Cryptology, 9(1): 19–??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

//www.degruyter.com/view/
j/gcc.2017.9.issue-1/gcc-
2017-0002/gcc-2017-0002.xml.

Conder:2010:UHG

[Con10] Marston Conder. An update on Hurwitz groups. Groups. Complexity. Cryptology, 2(1):35–49, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Craven:2016:PEA

[CR16] Matthew J. Craven and Daniel Robertz. A parallel evolutionary approach to solving systems of equations in polycyclic groups. Groups. Complexity. Cryptology, 8(2):109–125, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Chum:2010:LSS

[CZ10] Chi Sing Chum and Xiaowen Zhang. The latin squares and the secret sharing schemes. Groups. Complexity. Cryptology, 2(2):175–202, 2010. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Das:2015:KEF

[Das15] Manik Lal Das. Key-escrow free multi-signature scheme using bilinear pairings. Groups. Complexity. Cryptology, 7(1):47–57, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Diekert:2012:CRC

[DDM12] Volker Diekert, Andrew Duncan, and Alexei G. Myasnikov. Cyclic rewriting and conjugacy problems. Groups. Complexity.

Cryptology, 4(2):321–355, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Distler:2015:GES

[DE15] Andreas Distler and Bettina Eick. Group extensions with special properties. Groups. Complexity. Cryptology, 7(1): 1–10, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Dean:2009:MPF

[Dea09] Margaret H. Dean. Metabelian product of a free nilpotent group with a free abelian group. Groups. Complexity. Cryptology, 1(2):169–180, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Duquesne:2016:MSC

[DG16] Sylvain Duquesne and Loubna Ghammam. Memory-saving computation of the pairing final exponentiation on BN curves. Groups. Complexity. Cryptology, 8(1):75–90, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Eick:2017:IPT

[EE17] Bettina Eick and Ann-Kristin Engel. The isomorphism problem for torsion free nilpotent groups of Hirsch length at most 5. Groups. Complexity. Cryptology, 9(1):55– ??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

//www.degruyter.com/view/

                                        j/gcc.2017.9.issue-1/gcc-
                                        2017-0004/gcc-2017-0004.xml.

Eftekhari:2012:DHK

[Eft12] Mohammad Eftekhari. A Diffie– Hellman key exchange protocol using matrices over noncommutative rings. Groups. Complexity. Cryptology, 4(1):167–176, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Eick:2016:AGF

[Eic16] Bettina Eick. The automorphism group of a finitely generated virtually abelian group. Groups. Complexity. Cryptology, 8(1):35–45, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Elder:2010:SGP

[ER10] Murray Elder and Andrew Rechnitzer. Some geodesic problems in groups. Groups. Complexity. Cryptology, 2(2): 223–229, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Elder:2012:CTG

[ERW12] Murray Elder, Andrew Rechnitzer, and Thomas Wong. On the cogrowth of Thompson’s group F. Groups. Complexity. Cryptology, 4(2):301–320, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Elder:2016:TGC

[ET16] Murray Elder and Jennifer Taback. Thompson’s group F is 1-counter graph automatic. Groups. Complexity. Cryptology, 8(1):21–33, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fouotsa:2016:FAP

[FC16] Emmanuel Fouotsa and Abdoul Aziz Ciss. Faster Ate pairing computation on Selmer’s model of elliptic curves. Groups. Complexity. Cryptology, 8(1): 55–67, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Falconer:2011:GRE

[FFK11] Kenneth J. Falconer, Benjamin Fine, and Delaram Kahrobaei. Growth rate of an endomorphism of a group. Groups. Complexity. Cryptology, 3(2): 285–300, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fine:2014:RSA

[FGRS14] Benjamin Fine, Anthony Gaglione, Gerhard Rosenberger, and Dennis Spellman. Reflections on some aspects of infinite groups. Groups. Complexity. Cryptology, 6(2):81–91, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fine:2018:OGE

[FGRS18] Benjamin Fine, Anthony Gaglione, Gerhard Rosenberger, and Dennis Spellman. Orderable groups, elementary theory, and the Kaplansky conjecture. Groups. Complexity. Cryptology, 10(1):

43–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL

                                        https://www.degruyter.com/
                                        view/j/gcc.2018.10.issue-
                                        1/gcc-2018-0005/gcc-2018-
                                        0005.xml.

Fine:2009:GSG

[FMR09] Benjamin Fine, Alexei Myasnikov, and Gerhard Rosenberger. Generic subgroups of group amalgams. Groups. Complexity. Cryptology, 1(1):51–61, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fine:2013:SSS

[FMR13] Benjamin Fine, Anja I. S. Moldenhauer, and Gerhard Rosenberger. A secret sharing scheme based on the closest vector theorem and a modification to a private key cryptosystem. Groups. Complexity. Cryptology, 5(2): 223–238, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fine:2011:NFR

[FR11] Benjamin Fine and Gerhard Rosenberger. A note on faithful representations of limit groups. Groups. Complexity. Cryptology, 3(2):349–355, 2011. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Fine:2013:FRL

[FR13] Benjamin Fine and Gerhard Rosenberger. Faithful representations of limit groups II.

Groups. Complexity. Cryptology, 5(1):91–96, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Friedl:2017:EPG

[Fri17] Stefan Friedl. An elementary proof of the group law for elliptic curves. Groups. Complexity. Cryptology, 9(2):117–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigorchuk:2019:KAB

[GG19] Rostislav Grigorchuk and Dima Grigoriev. Key agreement based on automaton groups. Groups. Complexity. Cryptology, 11(2): 77–81, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigoriev:2009:CPK

[GHP09] Dima Grigoriev, Edward A.

Hirsch, and Konstantin Pervyshev. A complete public-key cryptosystem. Groups. Complexity. Cryptology, 1(1):1–12, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Gilman:2014:GSC

[Gil14] Robert H. Gilman. Generalized small cancellation presentations for automatic groups. Groups. Complexity. Cryptology, 6(2): 93–101, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Gryak:2016:SPG

[GK16] Jonathan Gryak and Delaram Kahrobaei. The status of polycyclic group-based cryptography: a survey and open problems. Groups. Complexity. Cryptology, 8(2):171–186, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Gilman:2018:GWW

[GKS18a] Robert H. Gilman, Robert P. Kropholler, and Saul Schleimer. Groups whose word problems are not semilinear. Groups. Complexity. Cryptology, 10(2): 53–??, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Gribov:2018:PPK

[GKS18b] Alexey Gribov, Delaram Kahrobaei, and Vladimir Shpilrain. Practical private-key fully homomorphic encryption in rings. Groups. Complexity. Cryptology, 10(1):17–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2018.10.issue-1/gcc-2018-
0006/gcc-2018-0006.xml.

Gaglione:2009:ALF

[GLS09] Anthony M. Gaglione, Seymour Lipschutz, and Dennis Spellman. Almost locally free groups and a theorem of Magnus: some questions. Groups. Complexity. Cryptology, 1(2):181–198, 2009.

CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Gaglione:2015:AER

[GLS15] Anthony M. Gaglione, Seymour Lipschutz, and Dennis Spellman. An application of elementary real analysis to a metabelian group admitting integral polynomial exponents. Groups. Complexity. Cryptology, 7(1):59–68, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ghaffari:2018:MSV

[GM18] Mohammad Hossein Ghaffari and Zohreh Mostaghim. More secure version of a Cayley hash function. Groups. Complexity. Cryptology, 10(1):29–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2018.10.issue-1/gcc-2018-
0002/gcc-2018-0002.xml.

Garreta:2017:RNG

[GMO17] Albert Garreta, Alexei Miasnikov, and Denis Ovchinnikov. Random nilpotent groups, polycyclic presentations, and Diophantine problems. Groups. Complexity. Cryptology, 9(2): 99–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Gilman:2011:REF

[GMR11] Robert H. Gilman, Alexei Myasnikov, and Vitaliĭ Roman’kov. Random equations in free groups. Groups. Complexity. Cryptology, 3(2):257–284, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigoriev:2012:CHI

[GN12] Dima Grigoriev and Sergey Nikolenko. Continuous hard-to-invert functions and biometric authentication. Groups. Complexity. Cryptology, 4(1):19–32, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

grosseRebel:2009:TAT

[gRHR09] Volkmar große Rebel, Miriam Hahn, and Gerhard Rosenberger. The Tits alternative for Tsaranov’s generalized tetrahedron groups. Groups. Complexity. Cryptology, 1(2):207–216, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigoriev:2009:AMC

[GS09] Dima Grigoriev and Vladimir Shpilrain. Authentication from matrix conjugation. Groups. Complexity. Cryptology, 1(2): 199–205, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigoriev:2012:NLA

[GS12] Dima Grigoriev and Vladimir Shpilrain. No-leak authentication by the Sherlock Holmes method. Groups. Complexity. Cryptology, 4(1):177–189, 2012.

CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Grigoriev:2013:SOW

[GS13] Dima Grigoriev and Vladimir Shpilrain. Secrecy without oneway functions. Groups. Complexity. Cryptology, 5(1):31–52, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Hart:2015:FSA

[HHMHM15] Sarah Hart, Ivo Hedtke, Matthias Müller-Hannemann, and Sandeep Murthy. A fast search algorithm for ⟨m, m, m⟩ triple product property triples and an application for 5 × 5 matrix multiplication. Groups. Complexity. Cryptology, 7(1):31–46, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Habeeb:2013:DMR

[HK13] Maggie Habeeb and Delaram Kahrobaei. On the dimension of matrix representations of finitely generated torsion free nilpotent groups. Groups. Complexity. Cryptology, 5(2): 193–209, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Hedtke:2012:STA

[HM12] Ivo Hedtke and Sandeep Murthy.

Search and test algorithms for triple product property triples. Groups. Complexity. Cryptology, 4(1):111–133, 2012. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ho:2018:WPZ

[Ho18]  Meng-Che Ho.  The word

problem of z n is a multiple context-free language. Groups. Complexity. Cryptology, 10(1): 9–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL

                                        https://www.degruyter.com/
                                        view/j/gcc.2018.10.issue-
                                        1/gcc-2018-0003/gcc-2018-
                                        0003.xml.

Holt:2013:SAG

[HR13] Derek F. Holt and Sarah Rees. Shortlex automaticity and geodesic regularity in Artin groups. Groups. Complexity. Cryptology, 5(1):1–23, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ibrahim:2012:CMC

[Ibr12]  Fedaa Ibrahim.  Cylinders,

multi-cylinders and the induced action of Aut(F_{n}). Groups. Complexity. Cryptology, 4(2): 357–375, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ito:2010:FTT

[Ito10] Tetsuya Ito. On finite Thurstontype orderings of braid groups. Groups. Complexity. Cryptology, 2(2):123–155, 2010. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ivachev:2015:TDM

[Iva15] Artyom S. Ivachev. On transitive differentiable modulo p^{n} functions. Groups. Complexity. Cryptology, 7(2):183–190, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Jovanovic:2010:AAU

[JK10] Philipp Jovanovic and Martin Kreuzer. Algebraic attacks using SAT-solvers. Groups. Complexity. Cryptology, 2(2): 247–259, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kahrobaei:2009:DSN

[KA09] Delaram Kahrobaei and Michael Anshel. Decision and search in non-abelian Cramer–Shoup public key cryptosystem. Groups. Complexity. Cryptology, 1(2): 217–225, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kassabov:2010:PMR

[Kas10] Martin Kassabov. Presentations of matrix rings. Groups. Complexity. Cryptology, 2(1):51–57, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kahrobaei:2012:NCD

[KK12] Delaram Kahrobaei and Charalambos Koupparis. Noncommutative digital signatures. Groups. Complexity. Cryptology, 4(2):377–384, 2012. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kahrobaei:2013:PKE

[KKS13] Delaram Kahrobaei, Charalambos Koupparis, and Vladimir Shpilrain. Public key exchange using matrices over group rings. Groups. Complexity. Cryptology, 5(1):97–115, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kalka:2009:NSC

[KLT09] Arkadius Kalka, Eran Liberman, and Mina Teicher. A note on the shifted conjugacy problem in braid groups. Groups. Complexity. Cryptology, 1(2): 227–230, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kalka:2010:SCP

[KLT10] Arkadius Kalka, Eran Liberman, and Mina Teicher. Subgroup conjugacy problem for Garside subgroups of Garside groups. Groups. Complexity. Cryptology, 2(2):157–174, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Koblitz:2013:ALN

[KM13] Neal Koblitz and Alfred Menezes. Another look at non-uniformity. Groups. Complexity. Cryptology, 5(2):117–139, 2013. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kahrobaei:2019:SAA

[KMK19] Delaram Kahrobaei and Keivan Mallahi-Karai. Some applications of arithmetic groups in cryptography. Groups. Complexity. Cryptology, 11(1): 25–??, May 2019. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL

https://www.degruyter.com/
view/j/gcc.2019.11.issue-
1/gcc-2019-2002/gcc-2019-
2002.xml.

Kharlampovich:2014:IWU

[KMS14] Olga Kharlampovich, Alexei Myasnikov, and Denis Serbin. Infinite words and universal free actions. Groups. Complexity. Cryptology, 6(1):55–69, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kharlampovich:2019:ECC

[KMT19] Olga Kharlampovich, Alexei Myasnikov, and Alexander Taam. Effective construction of covers of canonical Homdiagrams for equations over torsion-free hyperbolic groups. Groups. Complexity. Cryptology, 11(2):83–101, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Kaluba:2018:CNE

[KN18] Marek Kaluba and Piotr W.

Nowak. Certifying numerical estimates of spectral gaps. Groups. Complexity. Cryptology, 10(1):33–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2018.10.issue-1/gcc-2018-
0004/gcc-2018-0004.xml.

Kappe:2016:CNS

[KNPS16] Luise-Charlotte Kappe, Daniela Nikolova-Popova, and Eric Swartz. On the covering number of small symmetric groups and some sporadic simple groups. Groups. Complexity. Cryptology, 8(2):135–154, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kotov:2018:ASS

[KPU18] Matvei Kotov, Dmitry Panteleev, and Alexander Ushakov. Analysis of secret sharing schemes based on Nielsen transformations. Groups. Complexity. Cryptology, 10(1):1–??, May 2018. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

                                        tronic). URL https://www.
                                        degruyter.com/view/j/gcc.
                                        2018.10.issue-1/gcc-2018-
                                        0001/gcc-2018-0001.xml.

Krammer:2013:AGA

[Kra13] Daan Krammer. An asymmetric generalisation of Artin monoids. Groups. Complexity. Cryptology, 5(2):141–167, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kreuzer:2009:AAG

[Kre09] Martin Kreuzer. Algebraic attacks galore! Groups. Complexity. Cryptology, 1(2):231–259, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kron:2010:CGR

[Krö10] Bernhard Krön. Cutting up graphs revisited — a short proof of Stallings’ structure theorem. Groups. Complexity. Cryptology, 2(2):213–221, 2010. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Kropholler:2019:CPC

[KS19] Robert P. Kropholler and Davide Spriano. Closure properties in the class of multiple contextfree groups. Groups. Complexity. Cryptology, 11(1):1–??, May 2019. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2019.11.issue-1/gcc-2019-
2004/gcc-2019-2004.xml.

Kalka:2013:NAK

[KT13] Arkadius Kalka and Mina Teicher. Non-associative key establishment for left distributive systems. Groups. Complexity. Cryptology, 5(2):169–191, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Lohrey:2012:ASC

[Loh12] Markus Lohrey. Algorithms on SLP-compressed strings: a survey. Groups. Complexity. Cryptology, 4(2):241–299, 2012. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Lakin:2009:SCW

[LT09] Stephen R. Lakin and Richard M.

Thomas. Space complexity and word problems of groups. Groups. Complexity. Cryptology, 1(2):261–273, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Longrigg:2009:PAC

[LU09] Jonathan Longrigg and Alexander Ushakov. A practical attack on a certain braid group based shifted conjugacy authentication protocol. Groups. Complexity. Cryptology, 1(2):275–286, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Lewis:2012:IEF

[LW12] Mark L. Lewis and James B. Wilson. Isomorphism in expanding families of indistinguishable groups. Groups. Complexity. Cryptology, 4(1): 73–110, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Maclachlan:2009:ENE

[Mac09]  C. Maclachlan.  Existence

and non-existence of torsion in maximal arithmetic Fuchsian groups. Groups. Complexity. Cryptology, 1(2):287–295, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Mahalanobis:2010:DLP

[Mah10] Ayan Mahalanobis. The discrete logarithm problem in the group of non-singular circulant matrices. Groups. Complexity. Cryptology, 2(1):83–89, 2010.

CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Miller:2011:ICM

[Mil11] Russell Miller. An introduction to computable model theory on groups and fields. Groups. Complexity. Cryptology, 3(1):25–45, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Miller:2014:FR

[Mil14] Charles F. Miller, III. Friends and relatives of BS(1, 2). Groups. Complexity. Cryptology, 6(2):73–80, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Monico:2015:CSU

[MN15] Chris Monico and Mara D.

Neusel. Cryptanalysis of a system using matrices over group rings. Groups. Complexity. Cryptology, 7(2):175–182, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Monico:2019:CHF

[Mon19] Chris Monico. Cryptanalysis of a hash function, and the modular subset sum problem. Groups. Complexity. Cryptology, 11(1):17–??, May 2019. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (elec-

tronic). URL https://www.
degruyter.com/view/j/gcc.
2019.11.issue-1/gcc-2019-
2001/gcc-2019-2001.xml.

Myasnikov:2014:DCF

[MR14] Alexei Myasnikov and Vitalii Roman’kov. Diophantine cryptography in free metabelian groups: theoretical base. Groups. Complexity. Cryptology, 6(2):103–120, 2014. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Myasnikov:2015:LDA

[MR15] Alexei Myasnikov and Vitaliĭ Roman’kov. A linear decomposition attack. Groups. Complexity. Cryptology, 7(1):81–94, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Mishchenko:2017:KPN

[MT17] Alexei Mishchenko and Alexander Treier. Knapsack problem for nilpotent groups. Groups. Complexity. Cryptology, 9(1): 87–??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

                                        //www.degruyter.com/view/
                                        j/gcc.2017.9.issue-1/gcc-
                                        2017-0006/gcc-2017-0006.xml.

Myasnikov:2009:CAA

[MU09] Alex D. Myasnikov and Alexander Ushakov. Cryptanalysis of the Anshel–Anshel–Goldfeld– Lemieux key agreement protocol. Groups. Complexity. Cryptology, 1(1):63–75, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Mosina:2011:SLL

[MU11a] Natalia Mosina and Alexander Ushakov. Strong law of large numbers on graphs and groups. Groups. Complexity. Cryptology, 3(1):67–103, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Myasnikov:2011:RVK

[MU11b] Alexei Myasnikov and Alexander Ushakov. Random van Kampen diagrams and algorithmic problems in groups. Groups. Complexity. Cryptology, 3(1):121–185, 2011. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Myasnikov:2014:QAD

[MU14] Alexey D. Myasnikov and Alexander Ushakov. Quantum algorithm for discrete logarithm problem for matrices over finite group rings. Groups. Complexity. Cryptology, 6(1):31–36, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Miasnikov:2017:LSC

[MV17] Alexei Miasnikov and Svetla Vassileva. Log-space conjugacy problem in the Grigorchuk group. Groups. Complexity. Cryptology, 9(1):77– ??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

//www.degruyter.com/view/
j/gcc.2017.9.issue-1/gcc-
2017-0005/gcc-2017-0005.xml.

Myasnikov:2009:GCC

[Mya09] Alex D. Myasnikov. Generic case complexity and one-way functions. Groups. Complexity. Cryptology, 1(1):13–31, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Majewicz:2009:PCN

[MZ09] Stephen Majewicz and Marcos Zyman. Power-commutative nilpotent R-powered groups. Groups. Complexity. Cryptology, 1(2):297–309, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Marko:2017:PKC

[MZJ17] František Marko, Alexandr N. Zubkov, and Martin Juráš. Public-key cryptosystem based on invariants of diagonalizable groups. Groups. Complexity. Cryptology, 9(1):31– ??, May 2017. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL https:

                                        //www.degruyter.com/view/
                                        j/gcc.2017.9.issue-1/gcc-
                                        2017-0003/gcc-2017-0003.xml.

Neumann-Brosig:2010:NHH

[NBR10] Matthias Neumann-Brosig and Gerhard Rosenberger. A note on the homology of hyperbolic groups. Groups. Complexity. Cryptology, 2(2):203–212, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Neupane:2012:TPK

[Neu12] Kashi Neupane. Two-party key establishment: from passive to active security without introducing new assumptions. Groups. Complexity. Cryptology, 4(1):1–17, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Noskov:2016:GCC

[NR16] Gennady A. Noskov and Alexander N. Rybalov. Generic case complexity of the graph isomorphism problem. Groups. Complexity. Cryptology, 8(1):9–20, 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Orevkov:2015:ARQ

[Ore15] Stepan Yu. Orevkov. Algorithmic recognition of quasipositive 4-braids of algebraic length three. Groups. Complexity. Cryptology, 7(2):157–173, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Osin:2009:UTT

[Osi09] D. Osin. On the universal theory of torsion and lacunary hyperbolic groups. Groups. Complexity. Cryptology, 1(2): 311–319, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Pramanik:2019:RSS

[PA19] Jyotirmoy Pramanik and Avishek Adhikari. Ramp secret sharing with cheater identification in presence of rushing cheaters. Groups. Complexity. Cryptology, 11(2):103–113, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Panteleev:2019:CSP

[PU19] Dmitry Panteleev and Alexander Ushakov. Conjugacy search problem and the Andrews– Curtis conjecture. Groups. Complexity. Cryptology, 11(1): 43–??, May 2019. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL

                                        https://www.degruyter.com/
                                        view/j/gcc.2019.11.issue-
                                        1/gcc-2019-2005/gcc-2019-
                                        2005.xml.

Pueschel:2016:HGD

[Pue16] Kristen Pueschel. Hydra group doubles are not residually finite. Groups. Complexity. Cryptology, 8(2):163–170, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Romankov:2012:EG

[Rom12] Vitaliĭ Roman’kov. Equations over groups. Groups. Complexity. Cryptology, 4(2):191–239, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Romankov:2015:NPP

[Rom15] Vitalii A. Roman’kov. New probabilistic public-key encryption based on the RSA cryptosystem. Groups. Complexity. Cryptology, 7(2):153–156, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Romankov:2016:NDA

[Rom16] Vitaliĭ Roman’kov. A nonlinear decomposition attack. Groups. Complexity. Cryptology, 8(2): 197–207, November 2016. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Romankov:2017:CCP

[Rom17] Vitaliĭ Roman’kov. Cryptanalysis of a combinatorial public key cryptosystem. Groups. Complexity. Cryptology, 9(2):125–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Romankov:2018:TGS

[Rom18] Vitaly Roman’kov. Two general schemes of algebraic cryptography. Groups. Complexity. Cryptology, 10(2):83–??, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Romanakov:2019:IVA

[Rom19] Vitaliĭ Roman’kov. An improved version of the AAG cryptographic protocol. Groups. Complexity. Cryptology, 11(1): 35–??, May 2019. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic). URL

https://www.degruyter.com/
view/j/gcc.2019.11.issue-
1/gcc-2019-2003/gcc-2019-
2003.xml.

Rosenberger:2009:SBB

[Ros09] Gerhard Rosenberger. A short biography of Ben Fine. Groups. Complexity. Cryptology, 1(2):

141–142, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Rosenmann:2013:ISF

[Ros13] Amnon Rosenmann. On the intersection of subgroups in free groups: echelon subgroups are inert. Groups. Complexity. Cryptology, 5(2):211–221, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Riley:2014:PWW

[RS14] Tim R. Riley and Andrew W. Sale. Palindromic width of wreath products, metabelian groups, and max-n solvable groups. Groups. Complexity. Cryptology, 6(2):121–132, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Rybalov:2013:GCD

[Ryb13] Alexander Rybalov. Generic complexity of the Diophantine problem. Groups. Complexity. Cryptology, 5(1):25–30, 2013. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Rybalov:2015:GCS

[Ryb15]  Alexander Rybalov.  On

the generic complexity of the searching graph isomorphism problem. Groups. Complexity. Cryptology, 7(2):191–193, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Rybalov:2017:GHB

[Ryb17] Alexander Rybalov. Generic hardness of the Boolean satisfiability problem. Groups. Complexity. Cryptology, 9(2):151–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Sahattchieve:2015:CHQ

[Sah15] Jordan Sahattchieve. On convex hulls and the quasiconvex subgroups of F_{m} × Z^{n}. Groups. Complexity. Cryptology, 7(1): 69–80, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Schridde:2016:CDL

[Sch16]  Christian Schridde.  Com-

puting discrete logarithms using O((log q)^{2}) operations from {+, −, ×, ÷, &}. Groups. Complexity. Cryptology, 8(2):91– 107, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shamsabad:2019:RNS

[SDR19] Mohammad Reza Mirzaee Shamsabad, Seyed Mojtaba Dehnavi, and Akbar Mahmoodi Rishakani. Randomized nonlinear software-oriented MDS diffusion layers. Groups. Complexity. Cryptology, 11(2):123–131, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Shevlyakov:2010:AGN

[She10] A. Shevlyakov. Algebraic geometry over natural numbers. The classification of coordinate monoids. Groups. Complexity. Cryptology, 2(1):91–111, 2010.

CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shevlyakov:2016:IAS

[She16] Artem N. Shevlyakov. On irreducible algebraic sets over linearly ordered semilattices. Groups. Complexity. Cryptology, 8(2):187–195, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shevlyakov:2019:GAU

[She19] Artem N. Shevlyakov. On group automorphisms in universal algebraic geometry. Groups. Complexity. Cryptology, 11(2): 115–121, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Shnaps:2009:WCP

[Shn09] Daniella Bak Shnaps. The word and conjugacy problem for shuffle groups. Groups. Complexity. Cryptology, 1(2):143–164, 2009. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shpilrain:2010:SWP

[Shp10] Vladimir Shpilrain. Search and witness problems in group theory. Groups. Complexity. Cryptology, 2(2):231–246, 2010. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shpilrain:2014:DBI

[Shp14] Vladimir Shpilrain. Decoybased information security. Groups. Complexity. Cryptology, 6(2):149–155, 2014. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shpilrain:2016:CLF

[SS16] Vladimir Shpilrain and Bianca Sosnovski. Compositions of linear functions and applications to hashing. Groups. Complexity. Cryptology, 8(2):155–161, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Shahryari:2017:DPV

[SS17] Mohammad Shahryari and Artem Shevlyakov. Direct products, varieties, and compactness conditions. Groups. Complexity. Cryptology, 9(2):159–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Seberry:2011:NGD

[STT11] Jennifer Seberry, Vinhbuu To, and Dongvu Tonien. A new generic digital signature algorithm. Groups. Complexity. Cryptology, 3(2):221–237, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Schleimer:2019:GTS

[SW19] Saul Schleimer and Bert Wiest.

Garside theory and subsurfaces: Some examples in braid groups. Groups. Complexity. Cryptology, 11(2):61–75, November 2019. ISSN 1867-1144 (print), 1869-6104 (electronic).

Shpilrain:2009:UDP

[SZ09] Vladimir Shpilrain and Gabriel Zapata. Using decision problems in public key cryptography. Groups. Complexity. Cryptology, 1(1):33–49, 2009. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Silva:2018:FGS

[SZ18] Pedro V. Silva and Alexander Zakharov. On finitely generated submonoids of virtually free groups. Groups. Complexity. Cryptology, 10(2):63–??, November 2018. ISSN 1867-1144 (print), 1869-6104 (electronic).

Timoshenko:2017:RSE

[Tim17] Evgeny I. Timoshenko. A remark on spherical equations in free metabelian groups. Groups. Complexity. Cryptology, 9(2): 155–??, November 2017. ISSN 1867-1144 (print), 1869-6104 (electronic).

Taback:2015:TBL

[TY15] Jennifer Taback and Sharif Younes. Tree-based language complexity of Thompson’s group F. Groups. Complexity. Cryptology, 7(2):135– 152, 2015. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ushakov:2016:ACK

[Ush16] Alexander Ushakov. Authenticated commutator key agreement protocol. Groups. Complexity. Cryptology, 8(2):127– 133, November 2016. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Vassileva:2011:PTC

[Vas11] Svetla Vassileva. Polynomial time conjugacy in wreath products and free solvable groups. Groups. Complexity. Cryptology, 3(1):105–120, 2011. CO-DEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Ventura:2014:GTO

[Ven14] Enric Ventura. Group-theoretic orbit decidability. Groups. Complexity. Cryptology, 6(2): 133–148, 2014. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Wassink:2011:SRT

[Was11] Bronlyn Wassink. Subgroups of R. Thompson’s group F that are isomorphic to F. Groups. Complexity. Cryptology, 3(2): 239–256, 2011. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Weiss:2010:SGL

[Wei10] Uri Weiss. On Shephard groups with large triangles. Groups. Complexity. Cryptology, 2(1): 1–34, 2010. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).

Wilson:2012:EAA

[Wil12] James B. Wilson. Existence, algorithms, and asymptotics of direct product decompositions, I. Groups. Complexity. Cryptology, 4(1):33–72, 2012. CODEN ???? ISSN 1867-1144 (print), 1869-6104 (electronic).